OpenStudy (anonymous):

Explain how the law of universal gravitation works using any two of the following items: a person, a two pound block, a tree, a whale, a mountain, the Moon, or the Earth AND describe how distance plays a role in what you described.

OpenStudy (anonymous):

any two bodies in the universe attract each other with a force that is directly proportional to the product of their masses and inversely proportional to the square of the distance between them. (courtesy Wikipedia) i think i am not supposed to answer this question.. may be i am not meant to...

OpenStudy (aaronq):

So lets pick a person (70 kg) and the moon \(7.35*10^{22}\) kg The formula as stated above is \(F_g=G\dfrac{m_1*m_2}{d^2}\) both of the objects experience the same force, but because they weight differently, the acceleration towards each other is not the same.   As the distance increases, the force experienced (and thus the acceleration) rapidly decreases for both of the objects. For example, when distance, d, is doubled, the force decreases to quarter of the original. \(\dfrac{1}{d^2}\rightarrow \dfrac{1}{(2d)^2}=\dfrac{1}{4d^2}\)